For his daily training, an athlete runs around a 15 m radius track five times in 2.5 minutes. What is his average speed in m/s?


1
Expert's answer
2021-09-07T09:50:47-0400

Given:

"R=15\:\rm m"

"t=2.5\:\rm min=150\;\rm s"

"n=5"


The period of rotation

"T=\frac{t}{N}=\frac{150\:\rm s}{5}=30\: \rm s"

The average speed

"v=\frac{2\pi R}{T}=\frac{6.28*15\;\rm m}{30\:\rm s}=19\:\rm m/s"